# General Chen-Ricci inequalities for Riemannian submersions and Riemannian maps

### v1

@@ 4.3 Applications
In this subsection, we present the Chen-Ricci inequalities for Riemannian maps to generalized complex and generalized Sasakian space forms. We also obtain Chen-Ricci inequalities for Riemannian maps to real, complex, real Kähler, Sasakian, cosymplectic, Kenmotsu, and $C(\alpha)$-space forms. Furthermore, we derive the Chen-Ricci inequalities for invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant Riemannian maps.
In the following theorem, we derive the Chen-Ricci inequalities when the target spaces are generalized complex and generalized Sasakian space forms. Toward this, we use the general form established in Theorem 4.1.
@@ 2. $(ii)$
In addition, the equality cases are identical to those of Theorem 4.1.
@@
Note that in view of [44], the item $(i)$ of above theorem makes sense for non-constants $c_{1}$ and $c_{2}$ only when $\dim(M_{2})=m_{2}=4$. However, for constants $c_{1}$ and $c_{2}$ the values of $m_{2}$ could be greater than $4$. Therefore, for particular values of $c_{i}$s, we have the following inequalities.
Corollary 4.3.
Let $F$ be a Riemannian map that satisfies the hypotheses of Theorem 4.1.
  1. $(i)$
    If ${M_{2}}(c)$ is a real space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)c+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
  2. $(ii)$
    If ${M_{2}}(c)$ is a complex space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)\frac{c}{4}+\frac{3c}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
  3. $(iii)$
    If ${M_{2}}(c)$ is a real Kähler space form with dimension $4$, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)\frac{\left(c+3\alpha\right)}{4}+\frac{3\left(c-\alpha\right)}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
  4. $(iv)$
    If ${M_{2}}(c)$ is a Sasakian space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{S},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{S}-\left(1+T_{\eta}\right)\frac{(c-1)}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $T_{S}=\left(r-1\right)\frac{(c+3)}{4}+\frac{3(c-1)}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
  5. $(v)$
    If ${M_{2}}(c)$ is a Kenmotsu space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{K},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{K}-\left(1+T_{\eta}\right)\frac{(c+1)}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $T_{K}=\left(r-1\right)\frac{(c-3)}{4}+\frac{3(c+1)}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
  6. $(vi)$
    If ${M_{2}}(c)$ is a cosymplectic space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{C},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{C}-\left(1+T_{\eta}\right)\frac{c}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $T_{C}=\left(r-1\right)\frac{c}{4}+\frac{3c}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
  7. $(vii)$
    If ${M_{2}}(c)$ is a $C(\alpha)$ space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{C(\alpha)},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{C(\alpha)}-\left(1+T_{\eta}\right)\frac{(c-\alpha^{2})}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $T_{C(\alpha)}=\left(r-1\right)\frac{(c+3\alpha^{2})}{4}+\frac{3(c-\alpha^{2})}{4}\|PF_{\ast}X\|^{2}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
In addition, the equality cases are identical to those of Theorem 4.2.
Remark 4.4.
We observe that the case $(i)$ in the previous corollary is the same as the case [29,  Theorem  3.1].
Similar to [51, Chapter 6], we can extend the notions of invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant, Riemannian maps to a complex (resp. generalized Sasakian) space form. Since these notions impose the same conditions on the distribution $({\rm range~}F_{\ast}={\cal D}_{1}\oplus{\cal D}_{2})$ with $J$ (resp. $\phi$), we are directly taking into account these conditions to prove the following result. We also note that since the proof is similar to the proof of Theorem 4.2, we skip it.
Theorem 4.5.
Let $F$ be a Riemannian map that satisfies the hypotheses of Theorem 4.1.
  1. $(1)$
    Suppose $(M_{2}(c),g_{2},J)$ is a complex space form. If
    1. $(i)$
      $F$ is invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq T_{I}=\left(r-1\right)\frac{c}{4}+\frac{3c}{4}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
    2. $(ii)$
      $F$ is anti-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq T_{AI}=\left(r-1\right)\frac{c}{4}+\frac{1}{4}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
    3. $(iii)$
      $F$ is semi-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}T_{I},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ T_{AI},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
    4. $(iv)$
      $F$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq T_{AI}+\frac{3c}{4}\cos^{2}\theta.$
    5. $(v)$
      $F$ is semi-slant such that ${\cal D}_{1}$ is invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}T_{I},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ T_{AI}+\frac{3c}{4}\cos^{2}\theta,&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
    6. $(vi)$
      $F$ is hemi-slant such that ${\cal D}_{1}$ is anti-invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}T_{AI},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ T_{AI}+\frac{3c}{4}\cos^{2}\theta,&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
  2. $(2)$
    Suppose $\left({M_{2}}\left(c_{1},c_{2},c_{3}\right),\phi,\xi,\eta,g_{2}\right)$ is a generalized Sasakian space form. If
    1. $(i)$
      $F$ is invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{I},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    2. $(ii)$
      $F$ is anti-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{AI},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    3. $(iii)$
      $F$ is semi-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{I},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &T_{AI},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    4. $(iv)$
      $F$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    5. $(v)$
      $F$ is semi-slant such that ${\cal D}_{1}$ is invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{I},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &T_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    6. $(vi)$
      $F$ is hemi-slant such that ${\cal D}_{1}$ is anti-invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&T_{AI},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &T_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &T_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
In addition, the equality cases are similar to those of Theorem 4.1.
Remark 4.6.
We observe that the item $(ii)$ of case $(1)$ in the previous theorem becomes the same as [29,  Theorem  4.2].
4.4 General framework of improved Chen-Ricci inequalities
In this subsection, we present a general form of the improved Chen-Ricci inequality for Riemannian maps between Riemannian manifolds. Toward this, we have the following theorem.
Theorem 4.7.
  1. $(i)$
    $B^{{\cal H}}=0$, or
  2. $(ii)$
    $r=2$ and ( F_) ( h_1,h_1) = 3μV_r+1, ( F_) ( h_2,h_2) = μV_r+1, ( F_) ( h_1,h_2) =μV_r+2
for some function $\mu$ with respect to some orthonormal local frame field $\left\{h_{1},h_{2}\right\}$.
Proof.
At a point $p\in{M_{1}}$, let $\left\{h_{i}\right\}_{i=1}^{r}$ and $\{V_{r+1},\dots,V_{m_{2}}\}$ be orthonormal bases for $(\ker F_{\ast})^{\perp}$ and $({\rm range~}~F_{\ast})^{\perp}$, respectively, and
By (4.2), we get
Using (4.19) in (4.20), we derive
[…]
4.5 Applications
In this subsection, we present the improved Chen-Ricci inequalities for Riemannian maps to generalized complex and generalized Sasakian space forms. We also obtain improved Chen-Ricci inequalities for Riemannian maps to real, complex, real Kähler, Sasakian, cosymplectic, Kenmotsu, and $C(\alpha)$-space forms. Furthermore, we derive the improved Chen-Ricci inequalities for invariant, anti-invariant, semi-invariant, slant, semi-slant, and hemi-slant Riemannian maps.
In the following theorem, we derive the improved Chen-Ricci inequalities when the target spaces are generalized complex and generalized Sasakian space forms. Toward this, we use the general form established in Theorem 4.7.
@@
In addition, the equality cases are identical to those of Theorem 4.7.
@@
Note that in view of [44], the item $(i)$ of above theorem makes sense for non-constants $c_{1}$ and $c_{2}$ only when $\dim(M_{2})=m_{2}=4$. However, for constants $c_{1}$ and $c_{2}$ the values of $m_{2}$ could be greater than $4$. Therefore, for particular values of $c_{i}$s, we have the following inequalities.
Corollary 4.9.
Let $F$ be a Riemannian map that satisfies the hypotheses of Theorem 4.7.
  1. $(i)$
    If ${M_{2}}(c)$ is a real space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)c+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
  2. $(ii)$
    If ${M_{2}}(c)$ is a complex space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)\frac{c}{4}+\frac{3c}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
  3. $(iii)$
    If ${M_{2}}(c)$ is a real Kähler space form with dimension $4$, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left(r-1\right)\frac{\left(c+3\alpha\right)}{4}+\frac{3\left(c-\alpha\right)}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
  4. $(iv)$
    If ${M_{2}}(c)$ is a Sasakian space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{S},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{S}-\left(1+T_{\eta}\right)\frac{(c-1)}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $IT_{S}=\left(r-1\right)\frac{(c+3)}{4}+\frac{3(c-1)}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
  5. $(v)$
    If ${M_{2}}(c)$ is a Kenmotsu space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{K},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{K}-\left(1+T_{\eta}\right)\frac{(c+1)}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $IT_{K}=\left(r-1\right)\frac{(c-3)}{4}+\frac{3(c+1)}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
  6. $(vi)$
    If ${M_{2}}(c)$ is a cosymplectic space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{C},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{C}-\left(1+T_{\eta}\right)\frac{c}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $IT_{C}=\left(r-1\right)\frac{c}{4}+\frac{3c}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
  7. $(vii)$
    If ${M_{2}}(c)$ is a $C(\alpha)$ space form, then
    ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{C(\alpha)},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{C(\alpha)}-\left(1+T_{\eta}\right)\frac{(c-\alpha^{2})}{4},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}),\end{array}\right$.
    where $IT_{C(\alpha)}=\left(r-1\right)\frac{(c+3\alpha^{2})}{4}+\frac{3(c-\alpha^{2})}{4}\|PF_{\ast}X\|^{2}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}$.
In addition, the equality cases are identical to those of Theorem 4.7.
Similarly to Theorem 4.5, taking into account the specific conditions imposed by the notions of Riemannian maps, we can also prove the following result. Since the proof is similar to the proof of Theorem 4.8, we skip it.
Theorem 4.10.
Let $F$ be a Riemannian map that satisfies the hypotheses of Theorem 4.7.
  1. $(1)$
    Suppose $(M_{2}(c),g_{2},J)$ is a complex space form. If
    1. $(i)$
      $F$ is invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq IT_{I}=\left(r-1\right)\frac{c}{4}+\frac{3c}{4}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
    2. $(ii)$
      $F$ is anti-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq IT_{AI}=\left(r-1\right)\frac{c}{4}+\frac{(r-1)}{4r}\left\|{\rm trace~}B^{{\cal H}}\right\|^{2}.$
    3. $(iii)$
      $F$ is semi-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}IT_{I},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ IT_{AI},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
    4. $(iv)$
      $F$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq IT_{AI}+\frac{3c}{4}\cos^{2}\theta.$
    5. $(v)$
      $F$ is semi-slant such that ${\cal D}_{1}$ is invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}IT_{I},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ IT_{AI}+\frac{3c}{4}\cos^{2}\theta,&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
    6. $(vi)$
      $F$ is hemi-slant such that ${\cal D}_{1}$ is anti-invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}IT_{AI},&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{1});\\ IT_{AI}+\frac{3c}{4}\cos^{2}\theta,&\text{if }~F_{\ast}X\in\Gamma({\cal D}_{2}).\end{array}\right$.
  2. $(2)$
    Suppose $\left({M_{2}}\left(c_{1},c_{2},c_{3}\right),\phi,\xi,\eta,g_{2}\right)$ is a generalized Sasakian space form. If
    1. $(i)$
      $F$ is invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{I},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    2. $(ii)$
      $F$ is anti-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{AI},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    3. $(iii)$
      $F$ is semi-invariant then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{I},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &IT_{AI},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    4. $(iv)$
      $F$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    5. $(v)$
      $F$ is semi-slant such that ${\cal D}_{1}$ is invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{I},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &IT_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
    6. $(vi)$
      $F$ is hemi-slant such that ${\cal D}_{1}$ is anti-invariant and ${\cal D}_{2}$ is slant with angle $\theta$ then,
      ${\rm Ric}^{{\cal H}}(X)\leq\left\{\begin{array}[]{ll}&IT_{AI},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{1}),\xi\in\Gamma({\rm range~}F_{\ast});\\ &IT_{AI}+3c_{2}\cos^{2}\theta,~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast})^{\perp};\\ &IT_{AI}+3c_{2}(1-\eta(F_{\ast}X)^{2})\cos^{2}\theta-\left(1+T_{\eta}\right)c_{3},~{\rm if}~F_{\ast}X\in\Gamma({\cal D}_{2}),\xi\in\Gamma({\rm range~}F_{\ast}).\end{array}\right$.
In addition, the equality cases are similar to those of Theorem 4.7.
Remark 4.11.
We observe that the item $(ii)$ of case $(1)$ in the previous theorem becomes the same as [29,  Theorem  4.9] when we take $J({\rm range~}F_{\ast})=(\rm range~F_{\ast})^{\perp}$.
5 Conclusion
We have obtained general forms of the Chen-Ricci inequalities for Riemannian submersions and Riemannian maps between Riemannian manifolds. These general forms result in new techniques that are easy, elegant, and fruitful in obtaining the Chen-Ricci inequalities with space forms. As applications of these general forms, we have obtained such inequalities with other space forms, such as real, complex, real Kähler, generalized complex, Sasakian, Kenmotsu, cosymplectic, $C(\alpha)$, generalized Sasakian space forms. Moreover, we have validated these inequalities with the verification of existing cases. In the future, they may be explored for other remaining space forms.

### v2

@@ 4.5 Applications
In this subsection, we obtain the Chen-Ricci and improved Chen-Ricci inequalities for Riemannian maps to generalized complex and generalized Sasakian space forms. Toward this, we use the general forms established in Theorems 4.1 and 4.2. Also, for any ${\cal Z}\in\Gamma(TM_{2})$, we decompose
$J{\cal Z}~\text{or}~\phi{\cal Z}=P{\cal Z}+Q{\cal Z},$ (4.16)
where $P{\cal Z}\in\Gamma({\rm range~}F_{\ast})$, $Q{\cal Z}\in\Gamma({\rm range~}F_{\ast})^{\perp}$ with
$\|PF_{\ast}h_{1}\|^{2}=\sum\limits_{j=2}^{r}\left(g_{2}(F_{\ast}h_{1},PF_{\ast}h_{j})\right)^{2}=\sum\limits_{j=1}^{r}\left(g_{2}(F_{\ast}h_{1},PF_{\ast}h_{j})\right)^{2}.$
First, we obtain the Chen-Ricci inequalities as follows.
@@ 2. $(ii)$
The equality conditions are identical to those of Theorem 4.1.
@@
Now, we obtain the improved Chen-Ricci inequalities as follows.
@@
The equality conditions are identical to those of Theorem 4.2.
@@
5 Conclusion with more applications and validations
In Theorems 3.1, 4.1 and 4.2, we have obtained general forms of the Chen-Ricci inequalities for Riemannian submersions and Riemannian maps. Further, by proving Theorems 3.3, 4.4 and 4.5, we showed that these general forms yield new techniques that are easy, elegant, and fruitful for obtaining Chen-Ricci inequalities with other space forms. In view of [25], the item $(i)$ of Theorem 3.3 makes sense for non-constants $c_{1}$ and $c_{2}$ only when $m_{1}=4$. Similarly, the item $(i)$ of Theorems 4.4 and 4.5 makes sense for non-constants $c_{1}$ and $c_{2}$ only when $\dim(M_{2})=m_{2}=4$. However, for constants $c_{1}$ and $c_{2}$, the corresponding values of $m_{1}$ or $m_{2}$ could be greater than $4$. We note that one can directly obtain Chen-Ricci inequalities with several special space forms by substituting the values of $c_{1}$, $c_{2}$, and $c_{3}$ from Table 2. Hence, Theorems 3.3, 4.4 and 4.5 are helpful to explore more particular cases, such as real, complex, real Kähler, Sasakian, Kenmotsu, cosymplectic, almost $C(\alpha)$ space forms.
From [33, 34, 35, 36, 27, 38], we have the notions of invariant, anti-invariant, Lagrangian, semi-invariant, slant, semi-slant, and hemi-slant Riemannian submersions from almost Hermitian manifolds onto Riemannian manifolds. In addition, from [21, 3, 14, 4, 37, 20], we have the notions of such structured Riemannian submersions from almost contact metric manifolds onto Riemannian manifolds. Similar to [32, Chapter 6], we can also extend the notions of such structured Riemannian maps to a complex or generalized Sasakian space form. These notions impose different-different conditions on the decomposition of the distributions $(\ker F_{\ast}={\cal D}_{1}\oplus{\cal D}_{2})$ and $({\rm range~}F_{\ast}={\cal D}_{1}\oplus{\cal D}_{2})$ with the complex structure $J$ or contact structure $\phi$. Using those conditions directly and putting the appropriate values of $c_{1}$, $c_{2}$ and $c_{3}$, one can also find Chen-Ricci inequalities for such structured Riemannian submersions and Riemannian maps. We observe that this approach successfully verifies the following existing inequalities.
Reference(s) Inequalities
[1] Equations (37), (38), and (62)
[2] Equations (44), (53), (66), (73), and (74)
[7] Theorems 6, 7, 8, 9, and 10
[8, 9] Theorems 4.1, 4.2, 4.3, 4.4, 4.5, and 4.6
[18] Theorem 4.1
[19] Theorems 4 and 5
[22] Theorems  3.1, 4.2, and 4.3
[28] Theorems 4.1, 4.2, 4.3, 4.6, 4.7, and 4.8
[30] Theorems 3.4, 3.5, and 3.6
Thus, the general forms obtained in this paper are powerful tools towards developing and verifying Chen-Ricci inequalities for Riemannian submersions and Riemannian maps. In the future, they may be explored for other remaining space forms.